# On Poincaré–Sobolev level involving fractional GJMS operators on hyperbolic space

### v1

@@
Abstract: This paper is devoted to a qualitative analysis of the Poincaré–Sobolev level associated with the fractional GJMS operators $\mathcal{P}_{s}$ $\bigl(s\in(0,\tfrac{n}{2})\setminus\mathbb{N}\bigr)$ on the hyperbolic space $\mathbb{H}^{n}$. In contrast to the integer-order case, when $s\notin\mathbb{N}$ the operator $\mathcal{P}_{s}$ does not enjoy the conformal covariance that allows one, in the upper half-space or ball model, to relate it to the Euclidean fractional Laplacian $(-\Delta)^{s}$; this link is crucial for importing Euclidean theory. We therefore introduce $\widetilde{\mathcal{P}}_{s}$ ($s>0$), which is conformally related to the $(-\Delta)^{s}$. Our purpose in the paper is to analyze the monotonicity, attainability, and strict-gap regions of the Poincaré–Sobolev levels associated with $\mathcal{P}_{s}$ and with $\widetilde{\mathcal{P}}_{s}$ with $s\in(0,\tfrac{n}{2})\setminus\mathbb{N}$.
@@ 1  Introduction and Main Results
which satisfies the conformal intertwining identities (1.3)–(1.4) for all $s>0$; see [24, Theorem 1.7]. This operator provides a convenient bridge between analysis on $\mathbb{H}^{n}$ and the Euclidean setting. Moreover, the precise relation between $\mathcal{P}_{s}$ and $\widetilde{\mathcal{P}}_{s}$ is given by [24, Corollary 5.3]:
@@ 1.1 Euclidean (Fractional) Laplacian and Integer-Order GJMS Operators
  • High dimensions relative to $s\in(0,1)$: $n\geq 4s$. One has ([3, Claim 14.1])
@@
    and consequently (1.15) admits a nontrivial weak solution for $\lambda\in(0,\lambda_{1,s}(\Omega)).$ ([3, Theorem 14.1])
  • Low dimensions relative to $s\in(0,1)$: $2s<n<4s$. There exists a constant $\lambda_{s}^{*}>0$ such that ([3, Proposition 16.4])
@@
    In this regime, the relation between $\lambda_{s}^{*}$ and $\lambda_{1,s}(\Omega)$ is unknown—even when $\Omega$ is the unit ball; consequently, the standard mountain–pass scheme alone does not guarantee existence. In [3, Theorem 16.1], the authors combine mountain–pass and linking arguments, treating separately the cases $\lambda_{s}^{*}<\lambda_{1,s}(\Omega)$ and $\lambda_{s}^{*}\geq\lambda_{1,s}(\Omega)$, and thereby obtain nontrivial solutions of (1.15) for every $\lambda>\lambda_{s}^{*}$ that is not a Dirichlet eigenvalue of $(-\Delta)^{s}$.
Building upon the aforementioned findings, we derive the following conclusion.
@@
    $\mu|\Omega|^{\frac{2s}{n}}\leq S_{n,s,\Omega}(\mu)<S_{n,s,\Omega}(\lambda_{1,s}\left(\Omega\right))=0\ \,{\rm for}\ \mu>\lambda_{1,s}(\Omega)\quad{\rm and}\quad S_{n,s,\Omega}(\lambda)=S_{n,s}\ \,{\rm for}\ \lambda\leq 0.$
@@
  • If $n\geq 4$, then for $\frac{n(n-2)}{4}<\lambda\leq\frac{(n-1)^{2}}{4},$ problem (1.16) admits a positive entire solution ([25, Theorem 1.5]); for $\lambda\leq\frac{n(n-1)}{4}$, problem (1.16) does not have any positive entire solution.
@@
  • If $n\geq 4$, then $0<\mu\leq\frac{1}{4}$ guarantees a positive entire solution to (1.17) [25, Theorem 1.5]; whereas for $\mu\;\leq\;0,$ problem (1.17) admits no positive entire solution [25, Theorem 1.6].
  • If $n=3$, then for $\mu\;\leq\;\frac{1}{4},$ problem (1.17) has no positive entire solution [25, Theorem 1.7].
@@
    When $n\geq 4k$, $\mathcal{G}_{n,k}[H_{n,k}]=(0,+\infty)$. $H_{n,k}(\lambda)$ is achieved and strictly decreasing for every $\lambda\in(0,\lambda_{0,k}^{\mathrm{conf}})$ when $k\geq 2$, while for $k=1,$ it is achieved if and only if $\lambda\in(0,\lambda_{0,1}^{\mathrm{conf}}]$ and strictly decreasing on $(0,\lambda_{0,k}^{\mathrm{conf}}]$, -∞=H_n,k(μ)¡0¡H_n,k(λ_0,k^conf)   for μ¿λ_0,k^conf and H_n,k(λ)=S_n,k for $\lambda\leq 0$.
@@
    When $2k+2\leq n\leq 4k-1$, there exists $\lambda_{k}^{\mathrm{conf}}\in(0,\lambda_{0,k}^{\mathrm{conf}})$ such that $(\lambda_{k}^{\mathrm{conf}},\infty)\subset\mathcal{G}_{n,k}[H_{n,k}]$. Moreover, $H_{n,k}(\lambda)$ is achieved and strictly decreasing for $\lambda\in(\lambda_{k}^{\mathrm{conf}},\lambda_{0,k}^{\mathrm{conf}})$, -∞=H_n,k(μ)¡0¡H_n,k(λ_0,k^conf)   for μ¿λ_0,k^conf and H_n,k(λ)=S_n,k for $\lambda\leq 0$.
@@ 1.2 Fractional GJMS Operators
For $s\in(0,\,n/4]\cap\mathbb{N}$, the above result recovers [23, Theorem 1.7] and yields an alternative proof thereof; whereas for $s\in(n/4,\,n/2)$, they established that $\lambda_{s}^{\mathrm{conf}}\in\bigl(0,\,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}\bigr).$
@@
$\lambda_{0,s}^{\mathrm{conf}}-b_{s}=\begin{cases}\displaystyle\frac{\Gamma\bigl(s+\tfrac{1}{2}\bigr)^{2}}{\pi},&\sin(\pi s)>0,\\[5.16663pt] \displaystyle\frac{1+\sin(\pi s)}{\pi}\,\Gamma\bigl(s+\tfrac{1}{2}\bigr)^{2},&\sin(\pi s)\leq 0.\end{cases}$ (1.21)
@@
    For $s\in\left(0,\frac{n}{4}\right]$, $\mathcal{G}_{n,s}[\widetilde{H}_{n,s}]=(0,+\infty)$. $\widetilde{H}_{n,s}(\lambda)$ is achieved and strictly decreasing for every $\lambda\in(0,\widetilde{\lambda}_{0,s}^{\mathrm{conf}})$, -∞=~H_n,s(μ)¡0¡~H_n,s(~λ_0,s^conf)   for μ¿~λ_0,s^conf and ~H_n,s(λ)=S_n,s for $\lambda\leq 0$.
@@
    For $s\in(\frac{n}{4},\frac{n-1}{2})$ when $n\geq 3$ and $s\in(\frac{1}{2},1)$ when $n=2$, then there exists $\widetilde{\lambda}_{s}^{\mathrm{conf}}\in(0,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}]$ such that $\bigl(\widetilde{\lambda}_{s}^{\mathrm{conf}},\infty\bigr)\;\subset\;\mathcal{G}_{n,s}\bigl[\widetilde{H}_{n,s}\bigr]$ and for any $\lambda\leq 0<\widetilde{\lambda}_{0,s}^{\mathrm{conf}}<\mu$ one has
@@
    For $s\in\left(0,\frac{n}{4}\right]\cap\mathcal{B}_{+}$, $(b_{s},\infty)\subset\mathcal{G}_{n,s}\bigl[H_{n,s}\bigr]$ and $H_{n,s}(\lambda)$ is achieved and strictly decreasing for every $\lambda\in(b_{s},\lambda_{0,s}^{\mathrm{conf}})$. Moreover, H_n,s(λ)=S_n,s  for $\lambda\leq 0$.
@@
Finally, $\mathcal{P}_{s}$ and $\widetilde{\mathcal{P}}_{s}$ are defined via spectral functional calculus. As a consequence, the most accessible information is encoded in the frequency side through explicit multipliers, whereas direct control in the physical variable is not immediate. An explicit kernel representation of $\widetilde{\mathcal{P}}_{s}$ must be involved for certain priori estimates, see Proposition 4.10.
@@ 2.1 Half-Space Model and Poincaré Ball Model
$\nabla_{\mathbb{H}^{n}}u(r,z)=r^{2}\big(\partial_{r}u(r,z)\,\partial_{r}+\nabla_{z}u(r,z)\big),$
@@
$d(x,y)=\log\frac{1+|T_{y}(x)|}{1-|T_{y}(x)|},\qquad x,y\in\mathbb{B}^{n}.$
@@ 2.2 Helgason Fourier Transform
$a(u,v):=\int_{\mathbb{B}^{n}}\langle\nabla u,\nabla v\rangle_{g}\,dV_{\mathbb{B}^{n}},\quad u,v\in C_{c}^{\infty}(\mathbb{B}^{n}),$
@@
$\mathcal{D}(a)=H^{1}(\mathbb{B}^{n}),$
@@
which coincides with Sobolev space $H^{2}(\mathbb{B}^{n})$, see [19, Appendix B].
@@
$h_{\beta,\theta}(x):=\Bigl(\frac{\sqrt{1-|x|^{2}}}{|x-\theta|}\Bigr)^{\,n-1-2\beta{\bf i}},\quad\forall\,x\in\mathbb{B}^{n},$
it satisfies
@@
$f(x)=\int_{\mathbb{R}}\int_{\mathbb{S}^{n-1}}\overline{h_{\beta,\theta}(x)}\,\hat{f}(\beta,\theta)\,\frac{d\sigma(\theta)\,d\beta}{|c(\beta)|^{2}},$ (2.4)
with convergence in $L^{2}(\mathbb{B}^{n})$.
@@
$\widehat{\Phi(-\Delta_{\mathbb{B}^{n}})f}(\beta,\theta)=\Phi\bigl(\lambda(\beta)\bigr)\,\hat{f}(\beta,\theta),\quad f\in\mathcal{D}\bigl(\Phi(-\Delta_{\mathbb{B}^{n}})\bigr).$
@@
The spectral theorem for multiplication operators on $L^{2}$ implies that the spectrum of $M_{\varphi}$ is the essential range of its symbol $\varphi$, see [19, Appendix A]. Since $\mathcal{H}$ is unitary, the spectrum is invariant under this unitary equivalence,
@@
$\widehat{(-\Delta_{\mathbb{B}^{n}})^{s}f}(\beta,\theta)=\bigl(\beta^{2}+\rho^{2}\bigr)^{s}\,\hat{f}(\beta,\theta),\quad\ f\in H^{2s}(\mathbb{B}^{n}).$
@@ 2.3 Fractional GJMS Operators and Inequality
For $\gamma\in\bigl(0,\frac{n}{2}\bigr)\setminus\mathbb{N}$, recall the conformal fractional GJMS operator $\mathcal{P}_{\gamma}$ on $\mathbb{H}^{n}$ has the explicit spectral representation (1.2). Equivalently, under the Helgason–Fourier transform, $P_{\gamma}$ acts as a spectral multiplier:
@@
However, $P_{\gamma}$ does not satisfy the following intertwining relations (1.3) (1.4). To recover a direct intertwining with the Euclidean fractional Laplacian under the conformal identifications of $\mathbb{H}^{n}$ with $\mathbb{R}^{n}_{+}$ and $\mathbb{B}^{n}$, Lu et al. introduced an auxiliary family of fractional operators $\widetilde{\mathcal{P}}_{\gamma}$ on $\mathbb{H}^{n}$ in [24, Theorem 1.7]. Equivalently, under the Helgason–Fourier transform one has the multiplier representation
@@
$\frac{\bigl|\Gamma(a+{\bf i}\lambda)\bigr|}{\bigl|\Gamma(b+{\bf i}\lambda)\bigr|}\geq\frac{\bigl|\Gamma(a)\bigr|}{\bigl|\Gamma(b)\bigr|},\qquad\text{provided that }\ (a+k)^{2}\geq(b+k)^{2},\ \forall k\in\mathbb{N}.$
Thus,
@@
Let $0<s<\tfrac{n}{2}$. Then the following Sobolev inequality holds:
@@
At the critical exponent $2_{s}^{*}=\frac{2n}{n-2s}$ one also has the conformal invariance of the $L^{2_{s}^{*}}$ norm:
@@
$\frac{\displaystyle\int_{\mathbb{B}^{n}}w_{\varepsilon}\,(-\Delta)^{s}w_{\varepsilon}\,dx}{\Bigl(\displaystyle\int_{\mathbb{B}^{n}}|w_{\varepsilon}|^{2_{s}^{*}}\,dx\Bigr)^{\!2/2_{s}^{*}}}\longrightarrow S_{n,s}\quad\varepsilon\rightarrow 0.$
@@
$\frac{\displaystyle\int_{\mathbb{B}^{n}}u_{\varepsilon}\,\widetilde{\mathcal{P}}_{s}u_{\varepsilon}\,dV_{\mathbb{B}^{n}}}{\Bigl(\displaystyle\int_{\mathbb{B}^{n}}|u_{\varepsilon}|^{2_{s}^{*}}\,dV_{\mathbb{B}^{n}}\Bigr)^{\!2/2_{s}^{*}}}=\frac{\displaystyle\int_{\mathbb{B}^{n}}w_{\varepsilon}\,(-\Delta)^{s}w_{\varepsilon}\,dx}{\Bigl(\displaystyle\int_{\mathbb{B}^{n}}|w_{\varepsilon}|^{2_{s}^{*}}\,dx\Bigr)^{\!2/2_{s}^{*}}}\longrightarrow S_{n,s}.$
@@
3  Analysis of Fractional Poincaré–Sobolev Levels

### v2

@@
Abstract: This paper is devoted to a qualitative analysis of the Poincaré–Sobolev level associated with the fractional GJMS operators $\mathcal{P}_{s}$ $\bigl(s\in(0,\tfrac{n}{2})\setminus\mathbb{N}\bigr)$ on the hyperbolic space $\mathbb{H}^{n}$. In contrast to the integer-order case, when $s\notin\mathbb{N}$ the operator $\mathcal{P}_{s}$ does not enjoy the conformal covariance that allows one, in the upper half-space or ball model, to relate it to the Euclidean fractional Laplacian $(-\Delta)^{s}$; this link is crucial for importing Euclidean theory. We therefore introduce $\widetilde{\mathcal{P}}_{s}$ ($s>0$), which is conformally related to the $(-\Delta)^{s}$. Our purpose in the paper is to analyze the monotonicity, attainability, and strict-gap regions of the Poincaré–Sobolev levels associated with $\mathcal{P}_{s}$ and $\widetilde{\mathcal{P}}_{s}$.
@@ 1  Introduction and Main Results
which satisfies the conformal intertwining identities (1.3)–(1.4) for all $s>0$; see [29, Theorem 1.7]. This operator provides a convenient bridge between analysis on $\mathbb{H}^{n}$ and the Euclidean setting. Moreover, the precise relation between $\mathcal{P}_{s}$ and $\widetilde{\mathcal{P}}_{s}$ is given by [29, Corollary 5.3]:
@@ 1.1 Euclidean (Fractional) Laplacian and Integer-Order GJMS Operators
Recently, research on the Brezis–Nirenberg problem has significantly expanded in scope and depth; for a comprehensive overview, we refer the reader to [24, 9, 10, 36, 37, 35] and the references therein.
@@
  • High dimensions relative to $s\in(0,1)$: $n\geq 4s$. One has ([33, Claim 14.1])
@@
    and consequently (1.15) admits a nontrivial weak solution for $\lambda\in(0,\lambda_{1,s}(\Omega)).$ ([33, Theorem 14.1])
  • Low dimensions relative to $s\in(0,1)$: $2s<n<4s$. There exists a constant $\lambda_{s}^{*}>0$ such that ([33, Proposition 16.4])
@@
    In this regime, the relation between $\lambda_{s}^{*}$ and $\lambda_{1,s}(\Omega)$ is unknown—even when $\Omega$ is the unit ball; consequently, the standard mountain–pass scheme alone does not guarantee existence. In [33, Theorem 16.1], the authors combine mountain–pass and linking arguments, treating separately the cases $\lambda_{s}^{*}<\lambda_{1,s}(\Omega)$ and $\lambda_{s}^{*}\geq\lambda_{1,s}(\Omega)$, and thereby obtain nontrivial solutions of (1.15) for every $\lambda>\lambda_{s}^{*}$ that is not a Dirichlet eigenvalue of $(-\Delta)^{s}$.
Recently, research on the Brezis–Nirenberg problem involving the nonlocal operators has been studied extensively, we refer the reader to [7, 39, 1, 34, 31, 17] and the references therein. Building upon the aforementioned findings, we derive the following conclusion.
@@
    $\lambda|\Omega|^{\frac{2s}{n}}\leq S_{n,s,\Omega}(\lambda)<S_{n,s,\Omega}(\lambda_{1,s}\left(\Omega\right))=0\ \,{\rm for}\ \lambda>\lambda_{1,s}(\Omega)\quad{\rm and}\quad S_{n,s,\Omega}(\lambda)=S_{n,s}\ \,{\rm for}\ \lambda\leq 0.$
@@
  • If $n\geq 4$, then for $\frac{n(n-2)}{4}<\lambda\leq\frac{(n-1)^{2}}{4},$ problem (1.16) admits a positive entire solution ([30, Theorem 1.5]); for $\lambda\leq\frac{n(n-1)}{4}$, problem (1.16) does not have any positive entire solution.
@@
  • If $n\geq 4$, then $0<\mu\leq\frac{1}{4}$ guarantees a positive entire solution to (1.17) [30, Theorem 1.5]; whereas for $\mu\;\leq\;0,$ problem (1.17) admits no positive entire solution [30, Theorem 1.6].
  • If $n=3$, then for $\mu\;\leq\;\frac{1}{4},$ problem (1.17) has no positive entire solution [30, Theorem 1.7].
@@
    When $n\geq 4k$, $\mathcal{G}_{n,k}[H_{n,k}]=(0,+\infty)$. $H_{n,k}(\lambda)$ is achieved and strictly decreasing for every $\lambda\in(0,\lambda_{0,k}^{\mathrm{conf}})$ when $k\geq 2$, while for $k=1,$ it is achieved if and only if $\lambda\in(0,\lambda_{0,1}^{\mathrm{conf}}]$ and strictly decreasing on $(0,\lambda_{0,k}^{\mathrm{conf}}]$, -∞=H_n,k(λ)¡0¡H_n,k(λ_0,k^conf)   for λ¿λ_0,k^conf   and   H_n,k(λ)=S_n,k for $\lambda\leq 0$.
@@
    When $2k+2\leq n\leq 4k-1$, there exists $\lambda_{k}^{\mathrm{conf}}\in(0,\lambda_{0,k}^{\mathrm{conf}})$ such that $(\lambda_{k}^{\mathrm{conf}},\infty)\subset\mathcal{G}_{n,k}[H_{n,k}]$. Moreover, $H_{n,k}(\lambda)$ is achieved and strictly decreasing for $\lambda\in(\lambda_{k}^{\mathrm{conf}},\lambda_{0,k}^{\mathrm{conf}})$, -∞=H_n,k(λ)¡0¡H_n,k(λ_0,k^conf)   for λ¿λ_0,k^conf   and   H_n,k(λ)=S_n,k for $\lambda\leq 0$.
@@ 1.2 Fractional GJMS Operators
For $s\in(0,\,n/4]\cap\mathbb{N}$, the above result recovers [28, Theorem 1.7] and yields an alternative proof thereof; whereas for $s\in(n/4,\,n/2)$, they established that $\lambda_{s}^{\mathrm{conf}}\in\bigl(0,\,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}\bigr).$
@@
$\lambda_{0,s}^{\mathrm{conf}}-b_{s}=\begin{cases}\displaystyle\frac{\Gamma\bigl(s+\tfrac{1}{2}\bigr)^{2}}{\pi},&\sin(\pi s)>0,\\[4.64996pt] \displaystyle\frac{1+\sin(\pi s)}{\pi}\,\Gamma\bigl(s+\tfrac{1}{2}\bigr)^{2},&\sin(\pi s)\leq 0.\end{cases}$ (1.21)
@@
    For $s\in\left(0,\frac{n}{4}\right]$, $\mathcal{G}_{n,s}[\widetilde{H}_{n,s}]=(0,+\infty)$, $\widetilde{H}_{n,s}(\lambda)$ is achieved and strictly decreasing for $\lambda\in(0,\widetilde{\lambda}_{0,s}^{\mathrm{conf}})$, -∞=~H_n,s(μ)¡0¡~H_n,s(~λ_0,s^conf)   for μ¿~λ_0,s^conf   and   ~H_n,s(λ)=S_n,s for $\lambda\leq 0$.
@@
    For $s\in(\frac{n}{4},\frac{n-1}{2})$ when $n\geq 3$ and $s\in(\frac{1}{2},1)$ when $n=2$, then there exists $\widetilde{\lambda}_{s}^{\mathrm{conf}}\in(0,\widetilde{\lambda}_{0,s}^{\mathrm{conf}}]$ such that $\bigl(\widetilde{\lambda}_{s}^{\mathrm{conf}},\infty\bigr)\;\subset\;\mathcal{G}_{n,s}\bigl[\widetilde{H}_{n,s}\bigr]$ and for any $\lambda\leq 0<\widetilde{\lambda}_{0,s}^{\mathrm{conf}}<\mu$,
@@
    For $s\in\left(0,\frac{n}{4}\right]\cap\mathcal{B}_{+}$, $(b_{s},\infty)\subset\mathcal{G}_{n,s}\bigl[H_{n,s}\bigr]$ and $H_{n,s}(\lambda)$ is achieved and strictly decreasing for $\lambda\in(b_{s},\lambda_{0,s}^{\mathrm{conf}})$. Moreover, H_n,s(λ)=S_n,s    for $\lambda\leq 0$.
@@
Finally, $\mathcal{P}_{s}$ and $\widetilde{\mathcal{P}}_{s}$ are defined via spectral functional calculus. As a consequence, the most accessible information is encoded in the frequency side through explicit multipliers, whereas direct control in the physical variable is not immediate. An explicit kernel representation must be involved for certain priori estimates, see Proposition 4.10.
@@ 2  Geometric and Analytic Preliminaries
In this section, we present two fundamental models of the hyperbolic space, which will be used interchangeably throughout the paper according to the needs of different arguments. We then recall the Helgason–Fourier transform on hyperbolic space and explain its connection with functional calculus, which allows us to define general spectral fractional Laplacian operators. In the final part, we introduce the fractional GJMS operators that constitute the main object of this work, and state a key inequality that will be used repeatedly in the sequel.
@@ 2.1 Half-Space Model and Poincaré Ball Model
$\nabla_{\mathbb{H}^{n}}u(r,z)=r^{2}\big(\partial_{r}u(r,z)\,\partial_{r}+\nabla_{z}u(r,z)\big)$
@@
$d(x,y)=\log\frac{1+|T_{y}(x)|}{1-|T_{y}(x)|}\quad\ {\rm for}\ \,x,y\in\mathbb{B}^{n}.$
@@ 2.2 Helgason Fourier Transform
${\bf a}(u,v):=\int_{\mathbb{B}^{n}}\langle\nabla u,\nabla v\rangle_{g}\,dV_{\mathbb{B}^{n}},\quad u,v\in C_{c}^{\infty}(\mathbb{B}^{n}),$
@@
$\mathcal{D}({\bf a})=H^{1}(\mathbb{B}^{n})\times H^{1}(\mathbb{B}^{n}),$
@@
which coincides with Sobolev space $H^{2}(\mathbb{B}^{n})$, see [23, Appendix B].
@@
$h_{\beta,\theta}(x):=\Bigl(\frac{\sqrt{1-|x|^{2}}}{|x-\theta|}\Bigr)^{\,n-1-2\beta{\bf i}},\quad\forall\,x\in\mathbb{B}^{n},$ (2.2)
then it satisfies
@@
$f(x)=\int_{\mathbb{R}}\int_{\mathbb{S}^{n-1}}\overline{h_{\beta,\theta}(x)}\,\hat{f}(\beta,\theta)\,\frac{d\sigma(\theta)\,d\beta}{|c(\beta)|^{2}}.$ (2.5)
@@
$\mathcal{F}\big(\Phi(-\Delta_{\mathbb{B}^{n}})f\big)(\beta,\theta)=\Phi\bigl(\lambda(\beta)\bigr)\,\hat{f}(\beta,\theta),\quad f\in\mathcal{D}\bigl(\Phi(-\Delta_{\mathbb{B}^{n}})\bigr).$
@@
The spectral theorem for multiplication operators on $L^{2}$ implies that the spectrum of $M_{\varphi}$ is the essential range of its symbol $\varphi$, see [23, Appendix A]. Since $\mathcal{H}$ is unitary, the spectrum is invariant under this unitary equivalence,
@@
$\mathcal{F}\big((-\Delta_{\mathbb{B}^{n}})^{k}f\big)(\beta,\theta)=\bigl(\beta^{2}+\rho^{2}\bigr)^{k}\,\hat{f}(\beta,\theta)\quad{\rm for}\ \,f\in H^{2k}(\mathbb{B}^{n}).$
@@ 2.3 Fractional GJMS Operators and Inequality
For $\gamma\in\bigl(0,\frac{n}{2}\bigr)\setminus\mathbb{N}$, recall the conformal fractional GJMS operator $\mathcal{P}_{\gamma}$ on $\mathbb{H}^{n}$ has the explicit spectral representation (1.2). Equivalently, under the Helgason–Fourier transform, $\mathcal{P}_{\gamma}$ acts as a spectral multiplier:
@@
However, $\mathcal{P}_{\gamma}$ does not satisfy the following intertwining relations (1.3) and (1.4). To recover a direct intertwining with the Euclidean fractional Laplacian under the conformal identifications of $\mathbb{H}^{n}$ with $\mathbb{R}^{n}_{+}$ and $\mathbb{B}^{n}$, Lu et al. introduced an auxiliary family of fractional operators $\widetilde{\mathcal{P}}_{\gamma}$ on $\mathbb{H}^{n}$ in [29, Theorem 1.7]. Equivalently, under the Helgason–Fourier transform one has the multiplier representation
@@
$\frac{\bigl|\Gamma(a+{\bf i}\lambda)\bigr|}{\bigl|\Gamma(b+{\bf i}\lambda)\bigr|}\geq\frac{\bigl|\Gamma(a)\bigr|}{\bigl|\Gamma(b)\bigr|}\quad\text{provided that }\ (a+k)^{2}\geq(b+k)^{2},\ \forall k\in\mathbb{N}$
and
@@
Let $0<s<\tfrac{n}{2}$, then the following Sobolev inequality holds:
@@
At the critical exponent $2_{s}^{*}=\frac{2n}{n-2s}$, one also has the conformal invariance of the $L^{2_{s}^{*}}$ norm:
@@
$\frac{\displaystyle\int_{\mathbb{B}^{n}}w_{\varepsilon}\,(-\Delta)^{s}w_{\varepsilon}\,dx}{\Bigl(\displaystyle\int_{\mathbb{B}^{n}}|w_{\varepsilon}|^{2_{s}^{*}}\,dx\Bigr)^{\!2/2_{s}^{*}}}\to S_{n,s}\quad{\rm as}\ \,\varepsilon\rightarrow 0^{+}.$
@@
$\frac{\displaystyle\int_{\mathbb{B}^{n}}u_{\varepsilon}\,\widetilde{\mathcal{P}}_{s}u_{\varepsilon}\,dV_{\mathbb{B}^{n}}}{\Bigl(\displaystyle\int_{\mathbb{B}^{n}}|u_{\varepsilon}|^{2_{s}^{*}}\,dV_{\mathbb{B}^{n}}\Bigr)^{\!2/2_{s}^{*}}}=\frac{\displaystyle\int_{\mathbb{B}^{n}}w_{\varepsilon}\,(-\Delta)^{s}w_{\varepsilon}\,dx}{\Bigl(\displaystyle\int_{\mathbb{B}^{n}}|w_{\varepsilon}|^{2_{s}^{*}}\,dx\Bigr)^{\!2/2_{s}^{*}}}\to S_{n,s}\quad{\rm as}\ \,\varepsilon\rightarrow 0^{+}.$
@@
3  Hardy Lower Bounds for Fractional Hardy–Sobolev–Maz’ya Inequalities
This section develops the key test-function estimates needed to analyze the fractional Poincaré–Sobolev levels associated with $\widetilde{\mathcal{P}}_{s}$ and ${\mathcal{P}}_{s}$. We introduce the standard bubble $U_{\varepsilon}$ and its cut-off version $w_{\varepsilon}$, and quantify precisely how the truncation affects the key quantities: the critical norm, the $L^{2}$-mass of the corresponding test function $u_{\varepsilon}$, and the fractional Dirichlet energy. These estimates provide the core test-function machinery for the sequel, and are the main input for the proofs of Theorem 1.3 and Theorem 1.4.